OpenStudy (anonymous):

how about this question given that y=2x+4/x-2 and dy/dx=k/(x-2)^2.find the value of k

OpenStudy (anonymous):

my typing sucks. let me try again \[f'(x)=5(3x+1)^4\times 3=15(3x+1)^4\] by the chain rule \[f'(-1)=15(-3+1)^4\] whatever that is

OpenStudy (anonymous):

like this the answer?

sam (.sam.):

Well, what's 15(-3+1)^4 ?

OpenStudy (anonymous):

it looks like the function being differentiated is a rational function (the quotient of two polynomials). If the function is actually (2x+4)/(x-2), then you would differentiate the rational function using the quotient rule. Direct comparison would then lead to a value for k.